You have acquired a new CT scanner at a cost of $750,000. You expect to perform 7,000 procedures per year over the estimated 5-year life of the scanner. Assuming no salvage value and an annual increase in replacement cost of 10 percent, what capital charge per procedure should the hospital levy to provide for replacement cost in the second year, Explain? (Ignore financing costs or investment income offsets)

Please show work as I want to understand how to solve....
Answer #1

The expected cost in the second year is

750,000*1.1*1.1= 907,500

Life is 5 years, so cost to be amortized over 5 years

So capital charge per procedure = 907,500/(5*7000) = 25.928